OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form.

OpenStudy (campbell_st):

slope intercept form is y = mx + b m = slope and b is the y intercept distribute the right hand side then add 4 to both sides of your equation

OpenStudy (anonymous):

im still not understanding!?

OpenStudy (campbell_st):

well you have to distribute... which gives y - 4 = -2x + 10 last step, add 4 to both sides of he equation... then you'll have the equation of the line in slope intercept form.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its 8y = 12

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

nope add the 4 y - 4= -2x + 10 + 4 4 ------------------ y = -2x + 14

OpenStudy (anonymous):

so y-8= -2x + 14

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

nope... have a look at my solution above... the task is to get a single y equal to a combination of x's and constants... you had after distributing y - 4 = -2x + 10 to move the -4 to the right hand side of the equation you add 4 to both sides... and -4 + 4 = 0 so the left hand side is left with a single y and the equation becomes y = -2x + 14

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks so much for your help! I think I got it! :D